Two crews working toward each other are 21.25 miles apart. One crew paves 1.4 miles of highway per day, and the other paves 1.1 miles per day. How long will it take them to meet? Give the answer to one decimal place, if necessary.
Answer by mananth(16949)

Crew I -------- 1.4 miles per day
Crew II -------- 1.1 miles per day

they are paving towards each other
so add up rate of work
1.4 + 1.1 = 2.5 miles per day
Distance = 21.25 miles
time = d/r
time= 21.25 / 2.5
Time = 8.5 days